OpenStudy (igster1112):

There are 8 courses in 16 inches for Roman brick: how many courses of Roman brick are there in a 4-foot wall?

OpenStudy (igster1112):

Be back later ;)

OpenStudy (vuriffy):

4*12 = 48 inches. 12 inches per 1 foot. 16*3 = 48 inches. 3 * 8 = ? Is your answer of courses in a 4-foot wall.

OpenStudy (igster1112):

24